The length of a rectangle is twice the width of the rectangle.
The length of a diagonal of the rectangle is 25 cm.
Work out the area of the rectangle.
Give your answer as an integer.

Answer by thechamp1011(19)

LETS X=WIDTH
THEREFORE, LENGTH= 2X
USING THE PYTHEGOREAN THEOREM: X^2+2X^2=25^2
X^2+4X^2=625
5X^2=625
X^2=625/5=125
X=11.18
THE ARE OF REACTANGLE IS: A=L*W
A=2X*X=2X^2
A=2(11.18)^2=249.98CM